# Displacement correlations in a two-dimensional colloidal liquid and their relationship with shear strain correlations

### v1

@@ Acknowledgements.
We express our cordial gratitude to Grzegorz Szamel, Erika Eiser, Peter Schall, Hisao Hayakawa, Norihiro Oyama, Kunimasa Miyazaki, Atsushi Ikeda, Takenobu Nakamura, Hajime Yoshino, Kang Kim, Takahiro Hatano, and So Kitsunezaki for fruitful discussions, insightful comments and valuable suggestions. The first author (Ooshida) also thank Eni Kume and Shiladitya Sengupta for informative discussions on the occasion of recent international conferences held in the difficult time of the Covid19 pandemic. In addition, the assistance of Susumu Goto as a coauthor of Ref. [20] is gratefully acknowledged, especially in preparation of Fig. 1. This work was supported by Grants-in-Aid for Scientific Research (Kakenhi) JP-15K05213, JP-18K03459 and JP-21K03404, JSPS (Japan).

### v2

@@ Acknowledgements.
We express our cordial gratitude to Grzegorz Szamel, Erika Eiser, Peter Schall, Hisao Hayakawa, Norihiro Oyama, Kunimasa Miyazaki, Atsushi Ikeda, Takenobu Nakamura, Hajime Yoshino, Kang Kim, Takahiro Hatano, and So Kitsunezaki for fruitful discussions, insightful comments and valuable suggestions. The first author (Ooshida) also thank Eni Kume and Shiladitya Sengupta for informative discussions on the occasion of recent international conferences held in the difficult time of the Covid19 pandemic. Encouraging reactions to our arXiv preprint from Laurence Noirez and Alessio Zaccone are highly appreciated. Last but not least, the assistance of Susumu Goto as a coauthor of Ref. [24] is gratefully acknowledged, especially in preparation of Fig. 1. This work was supported by Grants-in-Aid for Scientific Research (Kakenhi) JP-15K05213, JP-18K03459 and JP-21K03404, JSPS (Japan).